The heat exhausted to the heat sink by a heat engine is twice the mechanical work done. What is its efficiency?


1
Expert's answer
2022-01-10T09:07:52-0500

By definition, the efficiency of an engine "\eta" is equal to the ratio of the heat energy "Q_h>0" taken from the high temperature heat source and "W>0" mechanical work done in one cycle:


"\eta = \dfrac{W}{Q_h}\times 100\%"

According to the energy conservation law:


"Q_h = W+Q_s"

where "Q_s>0" is the the heat exhausted to the heat sink. Since by the task "Q_s = 2W", find:


"Q_h = W+2W = 3W\\
\eta = \dfrac{W}{3W}\times 100\% \approx 33\%"

Answer. 33%.